Circular No. 7583

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001U IN NGC 5442 M. Schwartz, Cottage Grove, OR; and W. D. Li, University of Califor-

nia at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 17.7) on an unfiltered image taken with the 0.5-m Tenagra III telescope on Feb. 15.5 UT (confirmed at mag ∼ 17.3 on an earlier KAIT image taken on Feb. 3.5). SN 2001U is located at α = 14^{h}04^{m}42^{s}.94, δ = −9^{o}42′29′′.7 (equinox 2000.0), which is 4′′.7 west and 18′′.7 north of the nucleus of NGC 5442. A KAIT image taken on 2000 July 2.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001R IN NGC 5172 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001R (cf. IAUC 7579), obtained by M. Calkins on Feb. 16.47 UT with the Mt. Hopkins 1.5m telescope (+ FAST spectrograph), shows a continuum with well-defined P-Cyg lines of hydrogen, sodium, and iron, consistent with the plateau phase of normal type-II supernovae after maximum light. Adopting the NED recession velocity of 4030 km/s for NGC 5172, the expansion velocity derived from the minimum of the Hβ line is ∼ 2500 km/s.”

SUPERNOVA 2001S IN UGC 5491 Matheson et al. also report that a spectrum of SN 2001S (cf. IAUC

7580), obtained by Calkins as above on Feb. 16.32 UT, shows it to be a type-II supernova near maximum light. The spectrum consists of a blue continuum with superposed P-Cyg lines of hydrogen and Ca II. Adopting the NED recession velocity of 9100 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is ∼ 9000 km/s.

(20000) 2000 WR_{106}

T. L. Farnham, University of Texas, writes: “R- and V -band CCD photometry of (20000), obtained at the McDonald Observatory 2.1-m telescope on Jan. 24–27, shows a single-peaked rotational lightcurve with a probable period of 3.17 hr, though 2.78 and 3.67 hr cannot be ruled out. The 0.5-mag amplitude suggests that one hemisphere reflects at least 60 percent more light than the other. Less likely (with diameter ∼ 900 km; cf. IAUC 7554), the lightcurve could be double-peaked with a 6.34-hr period, which would imply that (20000) has an axial ratio a/b ≥ 1.6.”

2001 February 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green